OpenStudy (anonymous):

A student bought a box of crayons and 5 reams of paper for $54. She bought 5 boxes of crayons and 3 reams of paper for $50. What is the cost of each box of crayons and each ream of paper?

OpenStudy (anonymous):

Well, you have to set up a system of equations, so first set your variables: C = crayons, P = Paper. then set up equations for the sentence: A student bought a box of crayons and 5 reams of paper for $54. So the equation is C + 5P = 54 And the second one for: She bought 5 boxes of crayons and 3 reams of paper for $50. 5C + 3P = 50. Now, using these two equations, solve the system by elimination or substitution.

OpenStudy (anonymous):

I tried but when I tried solving for P, I got like 10.909090909090.

OpenStudy (anonymous):

Hmm, well let's go through it shall we? C + 5P = 54 ----> C = -5P + 54 5C + 3P = 50. then we can do the following by subbing first into second 5(-5P + 54) + 3P = 50 And you get: -25P + 270 + 3p = 50 Then rearrange to get: -22P = -220 So it follows that: 22P = 220, then divide by 22 to get P = 10. Use P = 10 (which is cost of paper) in the first equation to get C (cost of crayons).

OpenStudy (anonymous):

THANK YOU SO MUCH! You helped me out alot! HAVE A WONDERFUL EVENING!

OpenStudy (anonymous):

You too, and your welcome!

OpenStudy (anonymous):

I like that you don't just give peopl the answer!

OpenStudy (anonymous):

*people

OpenStudy (anonymous):

Oh yea, don't worry about that. :p